12x squared + 5x-2=0
Answer by srinivas123(23)

12x squared +5x-2=0
12x squared+8x-3x-2 = 0
4x(3x+2)-1(3x+2)=0
(4x-1)(3x+2)=0
-----> x=1/4 or x=-2/3.